OpenStudy (anonymous):

Solve the following equation for y: -4x +5y -20=0

OpenStudy (anonymous):

We want to isolate y so First we have to move -4x and -20 to the other side. You know how to?

OpenStudy (anonymous):

Do we add +4x to both sides?

OpenStudy (anonymous):

Yes and add 20 to both sides.

OpenStudy (anonymous):

You should get 5y = 4x + 20

OpenStudy (anonymous):

It's not over yet. We want 5 to left y alone. What should we do?

OpenStudy (anonymous):

I'm still having trouble getting 5y = 4x +20

OpenStudy (anonymous):

Oh nevermind I got it.

OpenStudy (anonymous):

So after we would add 4x to both sides right?

OpenStudy (anonymous):

No, We don't have to do it again. We just want to get 5 to the other side. You know how to?

OpenStudy (anonymous):

No, can you help me please? :)

OpenStudy (anonymous):

All we have to do is divide both side by 5.\[\large{5y = 4x + 20}\]\[\large{\frac{5y}{5} = \frac{4x + 20}{5}}\]\[\large{y = \frac{4x + 20}{5}}\]

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Yes is that all?

OpenStudy (anonymous):

Yeah, that it.

OpenStudy (anonymous):

Do we have to plug in x?

OpenStudy (anonymous):

oh nevermind, it's asking for y.